OpenStudy (anonymous):

6x+5y = -5 3x+5y = 10

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

Subtract the second equation from the first. That will get rid of y and you can solve for x.

OpenStudy (anonymous):

x = 5 ?

OpenStudy (ranga):

6x+5y = -5 3x+5y = 10 Subtract -------------- 3x = -15 x = -5

OpenStudy (anonymous):

y = -25?

OpenStudy (ranga):

no. try again.

OpenStudy (anonymous):

-2?

OpenStudy (ranga):

6x+5y = -5 x = -5 6(-5) + 5y = -5 -30 + 5y = -5 add 30 ti both sides: 5y = -5 + 30 = 25 y = 5